In an experiment on photoelectric effect, the slope of the cut-off voltage versus frequency of incident light is found to be 4.12 × 10–15 V s. Calculate the value of Planck’s constant.

Grade:11

1 Answers

Vikas TU
3 years ago
given, slope of cut-off voltage versus frequency of incident light = 4.12 × 10^-15 V/s
we know, slope of graph = Plank's constant/charge on electron
or, slope = h/e
so, h = slope × e
= 4.12 × 10^-15 V/s × 1.6 × 10^-19 C
= 6.592 × 10^-34 Js
hence, value of planks constant is 6.592 × 10^-34 J.s